# Modules as effective nodes in coarse-grained networks of Kuramoto oscillators

### v1

@@ Abstract
Most real-world networks exhibit a significant degree of modularity. Understanding the effects of such topology on dynamical processes is pivotal for advances in social and natural sciences. In this work we consider the dynamics of Kuramoto oscillators on modular networks and propose a simple coarse-graining procedure where modules are replaced by effective single oscillators. The method is inspired by EEG measurements, where very large groups of neurons under each electrode are interpreted as single nodes in a correlation network. We expose the interplay between intra-module and inter-module coupling strengths in keeping the coarse-graining process meaningful and show that its accuracy depends on the degree of intra-module synchronization. We show that, when modules are well synchronized, the phase transition from asynchronous to synchronous motion in networks with 2 and 3 modules is very well described by their respective reduced systems, regardless of the network structure connecting the modules. Application of the method to real networks with small modularity coefficients, on the other hand, reveals that the approximation is not accurate, although it still allows for the computation of the critical coupling and the qualitative behavior of the order parameter if the inter-module coupling is large enough.
@@ I Introduction
The original assumption that each oscillator interacts with all the others is a simplifying approximation that fails in many real systems, such as neurons in the brain, that are grouped in well-defined regions [23, 24, 25, 26, 27], and fireflies [28, 29, 30], that only interact with close neighbors. The extension of the Kuramoto model to networks [31, 32, 33, 34, 35] describing the set of possible pairwise interactions, revealed that the topology of connections has a large effect on the synchronization properties of the system, leading, for example, to frequency [36] and explosive synchronization [37, 38, 39, 40].
Here we propose an approximation to the dynamics of modular networks where each module is replaced by a single effective node. We compute the Kuramoto order parameter and study the conditions for the phase transition from disordered to synchronized state to be well represented by the reduced system. We find that such an approximation can be very good for a range of internal and inter-module coupling strengths. Moreover, we find that, under these conditions, the results are largely independent of the network of connections between modules, that can be ignored and replaced by a single link between the nodes representing the modules. This approach can be viewed as an alternative to other techniques based on the renormalization group that may rely on hyperbolic spaces [41], Laplacian matrices [42, 43, 44] and machine learning [45], which have been used to describe large networks in terms of fewer supernodes that retain the basic properties of the original system. Although our method only works well for networks with well defined modules, application to real networks lead to interesting, although not always accurate, approximations to the full dynamics. As examples we show numerical simulations for Zachary’s Karate club social network (two modules) and the C. Elegans gap junctions neural network (divided into three, five and ten modules).
This paper is organized as follows: in section II we develop our coarse graining procedure for modular networks and in sections III and IV we apply it to the simple cases of two and three modules respectively, using synthetic networks. In section V we apply the coarse graining method to Zachary’s Karate club and C. Elegans gap junctions networks. Finally in section VI we discuss our findings.
@@ II Coarse graining for modular networks
Networks can be described by an adjacency matrix $A$ containing information about the coupling between pairs of nodes. Here we consider only undirected and binary matrices, where $A_{ij}=A_{ji}=1$ if nodes $i$ and $j$ interact and $A_{ij}=0$ otherwise. The extension of the Kuramoto model to networks is given by
@@
It is usual in network theory to define modules as groups of nodes that are more densely connected to each other than to the rest of the network. Modular structures can be constructed by either decreasing the number of connections between nodes of different modules or by distinguishing between inner-module and outer-module connection strengths.
Here we construct synthetic modular networks as follows: modules are indexed by $\sigma=1,2,\dots,s$ and contain $N_{\sigma}$ nodes, or oscillators, with $N=\sum_{\sigma=1}^{s}N_{\sigma}$. We call $A^{\sigma\sigma^{\prime}}$ the block of the adjacency matrix between modules $\sigma$ and $\sigma^{\prime}$. We take pairs of oscillators that belong to the same module to be connected, i.e, $A_{ij}^{\sigma\sigma}=1$ for all $i,j\in\sigma$, and to have connection strength $\lambda_{\sigma\sigma}$. Nodes belonging to different modules, on the other hand, are connected with probability $p$ and strength $\lambda_{\sigma\sigma^{\prime}}=\lambda_{\sigma^{\prime}\sigma}$. Therefore, for $p=0$ the network has $s$ disjoint and fully connected modules, whereas for $p=1$ the modular character is manifested only by the different connection strengths.
@@
The average velocity is a variable regarding the module. If we were able to write Eq. (5) solely in terms of average module variables we could simplify the dynamics to $s$ equations, drastically reducing the dimensionality of the problem. This can be achieved in the special case where each and every module has a large enough internal synchrony such that the phases of oscillators belonging to the same module are about equal. In other words, if $\lambda_{\sigma\sigma}$ is large enough for $\theta_{\sigma,i}\approx\langle\theta_{\sigma}\rangle$, then we can lose the symbol $\langle\,.\,\rangle$ and arrive at
@@
which is the Kuramoto model in a fully connected network, independent of the connection matrix $A_{ij}^{\sigma\sigma^{\prime}}$. Since the coefficients $\lambda_{\sigma\sigma^{\prime}}$ do not need to be all equal, this constitutes an asymmetric network: the strength of the interaction that module $\sigma$ has with module $\sigma^{\prime}$ is not necessarily the same as $\sigma^{\prime}$ has with $\sigma$.
@@
Eq. (6) shows that, under model (4), the dynamics of $s$ modules is qualitatively the same as that of $s$ oscillators if high synchrony within modules is satisfied. This simple observation summarizes the analytical basis for the coarse-graining process. In the next sections we will quantify the validity of this result by simulating model (4) for $s=2$ and 3 to compare with the dynamics of 2 and 3 oscillators, respectively. Only symmetric cases will be treated, i.e, $\lambda_{\sigma\sigma}\equiv\lambda_{\text{in}}$ $\forall\sigma$ and $\lambda_{\sigma\sigma^{\prime}}=\lambda_{\sigma^{\prime}\sigma}\equiv\lambda/s$ $\forall\sigma\neq\sigma^{\prime}$. We shall see that Eq. (6) alone does not reproduce the behavior of the order parameter of the full modular network. However, when weights related to the degree of internal synchrony are introduced to renormalize the contribution of the coarse grained nodes, very good agreement is observed.
III Two modules
@@
where $\omega\equiv\omega_{1}-\omega_{2}$ is the oscillator’s frequency offset. For $\lambda>\omega$, Eq. (7) has a stable fixed point at $\phi^{\ast}=\arcsin{(\omega/\lambda)}$ and the order parameter at the stationary solution can be written as
$r=\frac{1}{\sqrt{2}}\sqrt{1+\cos{\phi^{*}}}=\frac{1}{\sqrt{2}}\sqrt{1+\sqrt{1-\left(\frac{\omega}{\lambda}\right)^{2}}}.$ (8)
Note that when the system is not synchronized the order parameter can still be written as $r^{2}=(1+\cos{\phi})/2$, but now with a non stable phase offset. This allows us to write a mean and a deviation of $r$ for asynchronized regions. The results are $\text{E}[r]=2/\pi$ and $\text{Var}[r]=1/2-4/\pi^{2}$.
Now we consider a network with 2 modules. Since typical modules do not have the same size we will show simulations for the case where $N_{1}=200$ and $N_{2}=100$. The frequency distributions $g(\omega)$ are Gaussians of width $\Delta=1$ and centered at $\omega_{1}$ and $\omega_{2}=0$ respectively. In order to apply the coarse graining procedure we must first tune the coupling constants to guarantee that the oscillators in each module are sufficiently synchronized. We define the local order parameters
$z_{\sigma}=r_{\sigma}e^{i\psi_{\sigma}}=\frac{1}{N_{\sigma}}\sum_{j=1}^{N_{\sigma}}e^{i\theta_{\sigma,j}}$ (9)
and set minimum synchronization thresholds $q_{\sigma}$ for each module’s local order parameter. In the simulations we initially fixed $q_{\sigma}=0.9$ for both modules. This means that, for a given value of the inter-module coupling parameter $\lambda$ we need to find $\lambda_{in}$ such that $r_{\sigma}\geq q_{\sigma}$ for $\sigma=1,2$.
@@
Figure 1 (a) illustrates the procedure for $\lambda=2$ on a network where the inter-module connection probability was set to $p=0.5$. We started by simulating Eq. (4) with internal coupling $\lambda_{\text{in}}=1$ for a fixed time interval of $\Delta t=8$. This assures the system went through the transient and stabilized. If, after this interval, $r_{\sigma}<q_{\sigma}$, we increase $\lambda_{\text{in}}$ by $0.1$ and continue the simulation for another $\Delta t$. We repeat the process until $r_{1}$ (green) and $r_{2}$ (yellow) surpass the boundary of $q_{\sigma}$ synchrony during a whole $\Delta t$. This determines the minimum $\lambda_{\text{in}}$ for the validity of the coarse graining procedure for given $\lambda$ and $p$.
@@
Figure 1: (a): Numerical integration of Eq. (4) for $\lambda=2$. Green, yellow and purple curves are the order parameters $r_{1}$, $r_{2}$ and $r$ and dashed line represents $q_{\sigma}=0.9$ synchrony threshold (all referring to left $y$ axis). Black ladder function shows $\lambda_{\text{in}}$ values (referring to the right $y$ axis). (b)-(c): Final $\lambda_{\text{in}}$ and its respective $\langle r\rangle$ with one standard deviation $\sigma$ above and bellow in purple. In (b) magenta and turquoise lines marks the average of critical $\lambda$ and $\lambda_{\text{in}}$ over 3 simulations. In (c) triangles represent $\sigma>0.01$ and circles otherwise, green ribbon marks the 2 oscillators limits where the top green curve is Eq. (8) and bottom boundary is Eq. (12). Theoretical mean and standard deviation for the asynchronized region are shown with dashed green lines. Panel (d) shows the analogous of panel (c) for minimal modular synchrony of $q_{\sigma}=0.7$, lower ribbon boundary is given by Eq. (11) with $q_{\sigma}=0.7$. Simulation values: $N_{1}=200$, $N_{2}=100$ and $p=0.5$. Gaussian frequency distributions with $\Delta=1$ and $(\omega_{1},\omega_{2})=(1.5,0)$.
@@
To compare the dynamics of the modular system to that of two oscillators, however, requires information on global synchronization. The purple curves in Figures 1 (a) represents the global order parameter. After reaching minimum $\lambda_{\text{in}}$ for internal synchrony we let the system evolve for a new time interval of $\Delta t=20$. This allows us to calculate the average and standard deviation of $r$ and, consequently, classify the state of the full system as synchronized or not. Note the stable value of $r$ indicating the small standard deviation.
@@
Calculating the standard deviation for a set of values of $\lambda$, instead of just the one presented in Figure 1 (a), we get panel (c) showing the average order parameter $\langle r\rangle$ as well as one unit of standard deviation $\sigma$ after reaching minimum $\lambda_{\text{in}}$, both in purple. For small $\lambda$, $r$ fluctuates and $\sigma$ is large, whereas for large $\lambda$ the order parameter converges to a stationary value with small standard deviation. We define the phase transition to global synchrony as the lowest $\lambda$ such that $\sigma$ is lower than the threshold of $1\%$. In the figure we use triangles for non-synchronized states and dots for synchronized states.
Fitting the global synchronization of two modules by Eq. (8) puts and upper boundary for the fit, shown as the upper green curve of Figure 1 (c). This happens because we force modules to have $q_{\sigma}\leq r_{\sigma}\leq 1$. Given that we slowly increase $\lambda_{\text{in}}$, a lower boundary for $r(\lambda)$ would constitute a better fit for the synchrony. To derive such equation we modify Eq. (2) and calculate the global order parameter taking into account the weights $q_{\sigma}$, related to the internal synchrony of each module:
$z=re^{i\psi}=\frac{1}{s}\sum_{\sigma=1}^{s}q_{\sigma}e^{i\theta_{\sigma}}.$ (10)
The weights $q_{\sigma}$ compensate for the fact that the module is not in perfect synchrony. For two modules with weights $q_{1}$ and $q_{2}$ we find
$r=\frac{1}{2}\sqrt{q_{1}^{2}+q_{2}^{2}+2q_{1}q_{2}\cos{\phi}}.$ (11)
which should replace Eq. (8).
Plugging in the minimum synchrony $q_{\sigma}=0.9$ for both modules we obtain
$r=\frac{0.9}{\sqrt{2}}\sqrt{1+\sqrt{1-\left(\frac{\omega}{\lambda}\right)^{2}}}.$ (12)
This is the lower bound curve shown in green in Figure 1 (c). The green ribbon marks the region between the upper and lower boundaries.
Note that it is impossible for the modules to synchronize at values above the upper boundary because it is a barrier of perfect oscillators which our modules would only asymptomatically reach when $\lambda_{\text{in}}\rightarrow\infty$. This also means that the synchronization can fall bellow the lower boundary as we see for some points of Figure 1 (c). Nonetheless, Eq. (12) constitutes a good fit for the global synchrony of the system. Going the opposite way and loosening our condition on minimum module synchrony to $r_{\sigma}\geq q_{\sigma}=0.7$ we get Figure 1 (d). The simulated values fall, mostly, within the ribbon which is clearly bigger due to the lower boundary now being Eq. (11) with $q_{1}=q_{2}=0.7$. Even though we can theoretically define the green ribbon as a reasonable region for the predicted global synchronization, the looser we let modules be the bigger is the region and consequently less insightful it is.
@@
Recall that for each $\lambda$ we had to determine a minimum $\lambda_{\text{in}}$ that fulfills the condition $r_{\sigma}\geq 0.9$. These values are shown in Figure 1 (b). We obtain a tent shaped curve that creates two regions: above the curve the coarse graining process is valid, and bellow the curve, it is not. Global phase transition is marked by the magenta line, while turquoise indicates the respective $\lambda_{\text{in}}$ for this globally synchronized state. The onset of synchronization is marked by the intersection of the magenta and turquoise lines.
@@
Observe that starting from global asynchrony, it gets harder to satisfy $r_{\sigma}\geq 0.9$ as we increase $\lambda$ thus demanding larger $\lambda_{\text{in}}$, i.e, the strength of connections among modules disrupts the cohesion within the modules. On the other hand, after transitioning to global synchrony the parameter $\lambda$ helps bringing modules close together and consequently, demands less from the inner-module coupling strength.
The onset of synchronization indicates both minimum $\lambda_{\text{in}}$ for $r_{\sigma}\geq 0.9$ and $\lambda$ for global synchrony. Figure 2 shows such $(\lambda_{\text{in}},\lambda)$ pairs for 5 different values of frequency offsets $\omega$ and 5 different inter-module connection probabilities $p$. Panel (a) shows that, for large enough $\lambda_{\text{in}}$, the global phase transition happens at about $\lambda=\omega$ which is precisely when the phase transition of a 2 oscillator system takes place, in accordance with the coarse-graining process. More interesting, the transition point does not depend upon the number of connections between modules, as expected from Eq. (6). Panel (b) shows that immediately after the transition to the globally synchronized state, the minimum $\lambda_{\text{in}}$ that would assure sufficient synchrony for the coarse-graining process decreases non linearly with the amount of connections between the modules, exposing the importance of network structure in keeping inner-module cohesion.

### v2

@@ Abstract
Most real-world networks exhibit a significant degree of modularity. Understanding the effects of such topology on dynamical processes is pivotal for advances in social and natural sciences. In this work we consider the dynamics of Kuramoto oscillators on modular networks and propose a simple coarse-graining procedure where modules are replaced by effective single oscillators. The method is inspired by EEG measurements, where very large groups of neurons under each electrode are interpreted as single nodes in a correlation network. We expose the interplay between intra-module and inter-module coupling strengths in keeping the coarse-graining process meaningful. We show that, when modules are well synchronized, the phase transition from asynchronous to synchronous motion in networks with 2 and 3 modules is very well described by their respective reduced systems, regardless of the network structure connecting the modules. Applications of the method to real networks with small modularity coefficients reveals that the approximation is also very accurate if oscillators in each module are identical. The method reproduces global synchronization patterns despite the low synchronizability of some modules, possibly allowing for the inference of the mean synchrony of each module when individual dynamics are not known.
@@ I Introduction
The original assumption that each oscillator interacts with all the others is a simplifying approximation that fails in many real systems, such as neurons in the brain, that are grouped in well-defined regions [23, 24, 25, 26, 27], fireflies [28, 29, 30], that only interact with close neighbors, and traffic routing, that depends critically on synchronization between neighbors [31]. The extension of the Kuramoto model to networks [32, 33, 34, 35, 36] describing the set of possible pairwise interactions, revealed that the topology of connections has a large effect on the synchronization properties of the system, leading, for example, to frequency [37] and explosive synchronization [38, 39, 40, 41, 31]. Another important aspect of the Kuramoto dynamics is how network properties, such as heterogeneity and symmetries, affect the paths to synchronization [42], which might involve the formation of clusters of synchronized oscillators resembling modules [43].
Here we propose an approximation to the dynamics of modular networks where each module is replaced by a single effective node. We compute the Kuramoto order parameter and study the conditions for the phase transition from disordered to synchronized state to be well represented by the reduced system. We find that such an approximation can be very good for a range of internal and inter-module coupling strengths. Moreover, we find that, under these conditions, the results are largely independent of the network of connections between modules, that can be ignored and replaced by a single link between the nodes representing the modules. This approach can be viewed as an alternative to other techniques based on the renormalization group that may rely on hyperbolic spaces [44], Laplacian matrices [45, 46, 47] and machine learning [48], which have been used to describe large networks in terms of fewer supernodes that retain the basic properties of the original system. One of the advantages of our method is that it can be applied to finite networks. It requires, however, that the modules are well defined, in a way that will become clear in the next sections. We will show examples with synthetic networks and also numerical simulations for Zachary’s Karate club social network (two modules) and the C. Elegans gap junctions neural network (divided into three, five and ten modules).
This paper is organized as follows: in sections II and III we develop our coarse graining procedure for modular networks and in sections IV and V we apply it to the simple cases of two and three modules respectively, using synthetic networks. In section VI we apply the coarse graining method to Zachary’s Karate club and C. Elegans gap junctions networks. Finally in section VII we discuss our findings.
@@ II Coarse graining for modular networks
Networks can be described by an adjacency matrix $A$ containing information about the coupling between pairs of nodes. Here we consider only undirected networks, where $A_{ij}=A_{ji}=1$ if nodes $i$ and $j$ interact and $A_{ij}=0$ otherwise. The extension of the Kuramoto model to networks is given by
@@
It is usual in network theory to define modules as groups of nodes that are more densely connected to each other than to the rest of the network. Modular structures can be constructed by either decreasing the number of connections between nodes of different modules or by distinguishing between intra-module and inter-module connection strengths.
In order to test our coarse graining procedure we will first apply it to synthetic modular networks. These networks are construct as follows: modules are indexed by $\sigma=1,2,\dots,s$ and contain $N_{\sigma}$ nodes, or oscillators, with $N=\sum_{\sigma=1}^{s}N_{\sigma}$. We call $A^{\sigma\sigma^{\prime}}$ the block of the adjacency matrix between modules $\sigma$ and $\sigma^{\prime}$. Nodes belonging to different modules are connected with probability $p$ and strength $\lambda_{\sigma\sigma^{\prime}}=\lambda_{\sigma^{\prime}\sigma}$. Therefore, for $p=0$ the network has $s$ disjoint modules. Similarly, oscillators belonging to the same module are connected according to $A^{\sigma\sigma}$ with strength $\lambda_{\sigma\sigma}$. We will say more about the structure of $A^{\sigma\sigma}$ shortly.
@@
The average velocity is a variable regarding the module. If we were able to write Eq. (5) solely in terms of average module variables we could simplify the dynamics to $s$ equations, drastically reducing the dimensionality of the problem. This can be achieved in the special case where each and every module has a large enough internal synchrony such that the phases of oscillators belonging to the same module are about equal. In other words, if $\lambda_{\sigma\sigma}$ is large enough for $\theta_{\sigma,i}\approx\langle\theta_{\sigma}\rangle$, then we can lose the symbol $\langle\,.\,\rangle$ to arrive at
@@
which is the Kuramoto model in a fully connected network, independent of the inter-module adjacency matrix $A^{\sigma\sigma^{\prime}}$. Note that, in this coarse-graining limit, the intra-module connection terms in Eq. (5) vanish due to the coherent phase of its oscillators. Therefore, the topological structure of the adjacency matrix $A$ becomes irrelevant and could assume any form. Since the coefficients $\lambda_{\sigma\sigma^{\prime}}$ do not need to be all equal, this constitutes an asymmetric network: the strength of the interaction that module $\sigma$ has with module $\sigma^{\prime}$ is not necessarily the same as $\sigma^{\prime}$ has with $\sigma$.
@@
Eq. (6) shows that, under model (4), the dynamics of $s$ modules is qualitatively the same as that of $s$ oscillators if high synchrony within modules is satisfied. This simple observation summarizes the analytical basis for the coarse-graining process. In the next sections we will quantify the validity of this result by simulating model (4) for $s=2$ and 3 to compare with the dynamics of 2 and 3 oscillators, respectively. Only symmetric cases will be treated, i.e, $\lambda_{\sigma\sigma}\equiv\lambda_{\text{in}}$ $\forall\sigma$ and $\lambda_{\sigma\sigma^{\prime}}=\lambda_{\sigma^{\prime}\sigma}\equiv\lambda/s$ $\forall\sigma\neq\sigma^{\prime}$. We shall see that Eq. (6) alone does not reproduce the behavior of the order parameter of the full modular network. However, when weights related to the degree of internal synchrony and relative module size are introduced to renormalize the contribution of the coarse grained nodes, very good agreement is observed.
Finally, we note that Eq. (6) is also valid for arbitrary modular networks, not just the synthetic types we described, as long as the requirement of high synchrony within modules is satisfied. Numerical simulations in section VI, however, show that in real networks not all modules synchronize easily. Even in these cases, the coarse-graining procedure works very well with minor adjustments that take that into account this local lack of synchrony.
III Order parameter
In this section we propose an expression to calculate the order parameter of the coarse-grained network that matches, within our approximation, that of the original network. We will actually estimate upper and lower bounds for $r$ considering the limits that maintain the coarse-graining method valid. We start from Eq. (2), that represents the global order parameter for the full network, and re-write it as
$z=\frac{1}{N}\sum_{\sigma=1}^{s}\sum_{j=1}^{N_{\sigma}}e^{i\theta_{\sigma,j}}.$ (7)
Defining local order parameters for each module as
$z_{\sigma}=\frac{1}{N_{\sigma}}\sum_{j=1}^{N_{\sigma}}e^{i\theta_{\sigma,j}}=r_{\sigma}e^{i\theta_{\sigma}}$ (8)
and $\rho_{\sigma}=N_{\sigma}/N$ where $N=\sum N_{\sigma}$ we can write $z$ in the compact form
$z=\sum_{\sigma=1}^{s}\rho_{\sigma}z_{\sigma}.$ (9)
Multiplying this equation by its complex conjugate gives
$r^{2}=\sum_{\sigma=1}^{s}\sum_{\sigma^{\prime}=1}^{s}\rho_{\sigma}\rho_{\sigma^{\prime}}r_{\sigma}r_{\sigma^{\prime}}e^{i(\theta_{\sigma}-\theta_{\sigma^{\prime}})}.$ (10)
Rearranging the terms and defining the phase offset between modules $\sigma$ and $\sigma^{\prime}$ by $\phi_{\sigma\sigma^{\prime}}=\theta_{\sigma}-\theta_{\sigma^{\prime}}$ we arrive at the final expression for the global order parameter
$r^{2}=\sum_{\sigma=1}^{s}\rho_{\sigma}^{2}r_{\sigma}^{2}+2\sum_{\begin{subarray}{c}\sigma,\sigma^{\prime}=1\\ \sigma^{\prime}>\sigma\end{subarray}}^{s}\rho_{\sigma}\rho_{\sigma^{\prime}}r_{\sigma}r_{\sigma^{\prime}}\cos{\phi_{\sigma\sigma^{\prime}}}.$ (11)
For the coarse-grained network with $s$ nodes the order parameter is given by a similar expression, but with $r_{\sigma}=1$. Also, by Eq. (6), in the coarse-graining regime the offset $\phi_{\sigma\sigma^{\prime}}$ between modules is the same offset obtained for the system of $s$ oscillators. Therefore, the expression
$r_{+}^{2}=\sum_{\sigma=1}^{s}\rho_{\sigma}^{2}+2\sum_{\begin{subarray}{c}\sigma,\sigma^{\prime}=1\\ \sigma^{\prime}>\sigma\end{subarray}}^{s}\rho_{\sigma}\rho_{\sigma^{\prime}}\cos{\phi_{\sigma\sigma^{\prime}}}.$ (12)
defines an upper bound for the order parameter of the full network, representing a situation where all modules are in perfect internal synchrony and the system behaves exactly as $s$ oscillators, weighted by size of modules they represent. We can determine the offset analytically for the case of two modules; for three and more modules we appeal to numerical solutions.
In the next sections we will use $r_{\sigma}$, the inner modular synchrony, as control parameters of our coarse-graining procedure. If $r_{\sigma}<<1$, modules loose cohesion and the coarse-graining becomes less accurate. In order to avoid that, we shall assume that the intensity of the internal coupling $\lambda_{\text{in}}$ is large enough so that modules have a minimum internal synchrony $q_{\sigma}$. Replacing $r_{\sigma}\to q_{\sigma}$ gives an estimate for $r$ that takes into account both the relative module sizes and their level of synchronization:
$r_{-}^{2}=\sum_{\sigma=1}^{s}\rho_{\sigma}^{2}q_{\sigma}^{2}+2\sum_{\begin{subarray}{c}\sigma,\sigma^{\prime}=1\\ \sigma^{\prime}>\sigma\end{subarray}}^{s}\rho_{\sigma}\rho_{\sigma^{\prime}}q_{\sigma}q_{\sigma^{\prime}}\cos{\phi_{\sigma\sigma^{\prime}}}.$ (13)
If the coarse-graining process is successful, the order parameter $r$ for the full network should satisfy $r_{-}<r<r_{+}$ where $r_{\pm}$ are computed with the reduced network with $s$ nodes. However, as we shall see in the next section and in section VI, some modules might not reach the desired level of synchrony. In that case $q_{\sigma}$ should be replaced by the actual value of $r_{\sigma}$ reached by the module.
IV Two modules
@@
where $\omega\equiv\omega_{1}-\omega_{2}$ is the oscillator’s frequency offset. For $\lambda>\omega$, Eq. (14) has a stable fixed point at $\phi^{\ast}=\arcsin{(\omega/\lambda)}=\sqrt{1-(\omega/\lambda)^{2}}$. It is straightforward to derive from Eq. (2) that the order parameter takes the form $r^{2}=(1+\cos{\phi})/2$. This result is independent of the synchronization status of the system. So when the system is not synchronized we can write a mean and a deviation of $r$. The results are $\text{E}[r]=2/\pi$ and $\text{Var}[r]=1/2-4/\pi^{2}$. This closes our short discussion on two oscillators systems.
@@
Now we proceed to simulate two modules systems. Since typical modules do not have the same size we will show simulations for the case where $N_{1}=200$ and $N_{2}=100$. In our first numerical experiment we assume that both modules are fully connected, i.e., $A^{\sigma\sigma}_{ij}=1$ if $i\neq j$ for $\sigma=1,2$. The frequency distributions $g(\omega)$ are Gaussians of width $\Delta=1$ and centered at $\omega_{1}=1.5$ and $\omega_{2}=0$, for modules 1 and 2 respectively. The synchrony threshold for both modules is set to $q_{\sigma}=0.9$. Figure 1 (a) illustrates the procedure to reach $q_{\sigma}$ on a network where the inter-module connection probability was set to $p=0.1$ and $\lambda=2$ . We started by simulating Eq. (4) with internal coupling $\lambda_{\text{in}}=1$ for a fixed time interval of $\Delta t=40$. This assures the system went through the transient and stabilized. If, after this interval, $r_{\sigma}<q_{\sigma}$, we increase $\lambda_{\text{in}}$ by $0.1$ and continue the simulation for another $\Delta t$. We repeat the process until $r_{1}$ (green) and $r_{2}$ (yellow) surpass the boundary of $q_{\sigma}$ synchrony during a whole $\Delta t$. This determines the minimum $\lambda_{\text{in}}$ for the validity of the coarse graining procedure for given $\lambda$ and $p$.
@@
Figure 1: Simulation of a two-module network with $N_{1}=200$, $N_{2}=100$ and Gaussian frequency distributions with $\Delta=1$ and $(\omega_{1},\omega_{2})=(1.5,0)$. Panels (a)-(c) show results for fully connected modules, with $A_{ij}^{\sigma\sigma}=1$, and inter-module nodes connected with probability $p=0.1$. Panels (d)-(f) correspond to an Erdös-Rényi network, where every pair of nodes is connected with probability $p=0.1$, irrespective of their location in module 1 or 2. Panels (a) and (d) show the order parameters $r_{1}$, $r_{2}$ and $r$ as a function of the time in green, yellow and purple, respectively, for $\lambda=2$. The black ladder function shows how $\lambda_{\text{in}}$ was increased over time to achieve high inter-module synchronization, right $y$-axis. Panels (b) and (e) show the global order parameter as a function of inter-module coupling $\lambda$ with its standard deviation; triangles represent $\sigma>0.01$ and circles otherwise. Green solid curves represent Eqs. (12) and (13). Theoretical mean and standard deviation for the asynchronized region are shown with green dashed lines. Panels (c) and (f) show the value of internal coupling for each external coupling such that desired internal synchrony $r_{\sigma}\geq q_{\sigma}$ was achieved. Magenta and turquoise lines marks the average of critical couplings over 3 independent simulations. A set of such pairs for various $\omega$ and $p$ are shown in Figure 2. Simulations in panels (b) and (e) where performed with $\lambda_{\text{in}}$ at the critical value.
@@
The purple curve in Figures 1 (a) represents the global order parameter. After reaching minimum $\lambda_{\text{in}}$ for internal synchrony $q_{\sigma}$ we let the system evolve for a final time interval. This allows us to calculate the average and standard deviation of $r$ and, consequently, classify the state of the full system as synchronized or not. Note the stable value of $r$, indicating small standard deviation and synchrony.
@@
Calculating the order parameter $\langle r\rangle$ and standard deviation $\sigma_{r}$ for a set of values of $\lambda$, instead of just the one presented in Figure 1 (a), we get panel (b), showing $\langle r\rangle$ with one unit of standard deviation after reaching minimum $\lambda_{\text{in}}$ (purple symbol and shaded area). For small $\lambda$, $r$ fluctuates and $\sigma_{r}$ is large whereas for large $\lambda$ the order parameter converges to a stationary value with small standard deviation. We define the phase transition to global synchrony as the lowest $\lambda$ such that $\sigma_{r}$ is lower than the threshold of $0.01$. In the figure we use triangles for non-synchronized states and dots for synchronized states. The green curves shown in panel (b) are related to a two oscillators system. The left horizontal lines are the mean and variance of asynchronous states of two oscillators, as discussed in the beginning of this section. Note the similarity with the simulated system. The right green functions show the upper bound $r_{+}$, Eq.(12) and $r_{-}$, Eq.(13) for the case of two modules. It is impossible for the modules to synchronize at values above the upper boundary because it is a barrier of perfect oscillators which our modules would only asymptomatically reach in particular cases when $\lambda_{\text{in}}\rightarrow\infty$. Nonetheless, Eq.(13) for two modules
$r_{-}=\sqrt{\rho_{1}^{2}q_{1}^{2}+\rho_{2}^{2}q_{2}^{2}+2\rho_{1}\rho_{2}q_{1}q_{2}\sqrt{1-\left(\frac{\omega}{\lambda}\right)^{2}}}$ (15)
represents a good fit for the global synchrony of the system.
@@
Recall that for each $\lambda$ we had to determine a minimum $\lambda_{\text{in}}$ that fulfills the condition $r_{\sigma}\geq q_{\sigma}$. These values are shown in Figure 1 (c). We obtain a tent shaped curve that creates two regions: above the curve the coarse graining process is valid, and bellow the curve, it is not. Global phase transition is marked by the magenta line, while turquoise indicates the respective $\lambda_{\text{in}}$ for this globally synchronized state. The onset of synchronization is marked by the intersection of the magenta and turquoise lines.
@@
Observe that before the phase transition, to satisfy the condition $r_{\sigma}\geq q_{\sigma}$ requires larger $\lambda_{\text{in}}$ as we increase $\lambda$, i.e, the strength of connections among modules disrupts the cohesion within the modules. On the other hand, after transitioning to global synchrony the parameter $\lambda$ helps bringing modules close together and consequently, demands less from the inner-module coupling strength.
In the previous experiment, both modules were considered to be fully connected. In order to relax this condition and study how the density of connections affect the coarse-graining, we show in panels (d)-(f) similar results for random modules, where internal nodes are also connected with probability $p=0.1$. Since we are setting the connection probability $p$ to be the same for every pair of nodes, the network falls into the Erdös-Rényi class. Modules are distinguished only by different connection strengths, $\lambda_{\text{in}}$ or $\lambda$.